OpenStudy (anonymous):

he slope between the two points (-1, -1) and (8, 0)

OpenStudy (anonymous):

M=rise/run is the change in y over the change in x. For example: Find the slope of the points 4,3 and -2/-1. Step 1. Take the y values 3 and -1 and subtract them. 3-(-1) =4 Step 2: Take the x values 4, -2 and subtract them. 4-(-2)=6 Step 3. Take the two numbers and make into a fraction. 3-(-1)/4-(-2)= 4/6 Step 4. Simplify 4/6 Answer 2/3